OpenStudy (anonymous):

Sharon is the youngest in her family. Her sister Anne is 3 years older than she is. Her brother Tom is 5 years older than Anne. The sum of their ages is 62. How old is Sharon?

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

Sharon is 17 because 17 + 3 = 20 20 + 5 = 25 17 + 20 + 25 = 62

OpenStudy (anonymous):

Hope I helped! :)

OpenStudy (anonymous):

thx